OpenStudy (taylormarie213):

One number exceeds another by 26. The sum of the numbers is 64. what are the numbers?

OpenStudy (anonymous):

let numbers be x and x+26.Then (x+x+26)=64 which means 2x+26=64 subtract 26 from both sides.you get 2x=38.That means x=38/2=19 Answer is 19 and 19+26=45